A batter hits a fly ball which leaves the bat 0.96 m above the ground at an angle of 62 ∘ with an initial speed of 28 m/s heading toward centerfield. Ignore air resistance.

A)How far from home plate would the ball land if not caught?

B)The ball is caught by the centerfielder who, starting at a distance of 110 m from home plate, runs straight toward home plate at a constant speed and makes the catch at ground level. Find his speed.
Answer #1

A)

Using kinematic equation,

S = ut + 0.5 at^2

-0.96 = (28 sin(62))t - 0.5 x 9.8 x t^2

4.9 t^2 - 24.72t - 0.96 = 0

Time, t = 5.983 sec

Distance, d = 28 cos(62) x 5.083 = 66.82 m

B)

V = (110 - 66.82)/5.083 = 8.50 m/s

Comment in case any doubt please rate my answer...